OpenStudy (anonymous):

how to evaluate csc 15 degrees on calculator?

hartnn (hartnn):

can you evaluate sin 15 degrees ? then csc 15 = 1/sin 15

OpenStudy (anonymous):

Make sure your calculator is set to DEG (degrees). since csc 15 = 1/sin (15) Enter 1/sin(15) and your answer is there on your screen.

hartnn (hartnn):

rimshaa, not actually , csc is not the inverse function of sin its just the reciprocal, csc x= 1/sin x (not sin^-1 x)

OpenStudy (anonymous):

THANKS GUYS <3

hartnn (hartnn):

welcome ^_^